# A non-singular universe out of Hayward black hole

### v2

@@ III Conformal Diagram
The integrand in the expression of $v(T,X)$ is regular, whereas the one for $u(T,X)$ is not. However, it can be handled with methods similar to those presented in Ref. [24] (see the discussion around Eqs. 2.29-3.31). These null coordinates still need to be appropriately compactified. The construction of the corresponding global coordinate system $(\tilde{u},\tilde{v})$ (covering both the exterior and the interior) is performed in Appendix B.
Figure 2: Conformal diagram representing our (quantum mechanically) modified OS collapse scenario. The FRW geometry, enclosed by curves marked by $r=0$ and $r=r_{b}$, is surrounded by the static Hayward black hole geometry, with horizons $X=X_{-}$ and $X=X_{+}$. We have taken $M=15$, $a_{0}=5$, $r_{b}=(4\pi/3)^{-1/3}$ and $\kappa=0.1$ (in Planck units).
The conformal diagram in Fig. 2 represents the collapse of a dust ball forming a Hayward black hole. The surface of the collapsing ball, at $r=r_{b}$, starts from $X\rightarrow\infty$ and crosses the pair of horizons $X^{+}$ and $X^{-}$, where $u\rightarrow\infty$ and $u\rightarrow-\infty$, respectively. Finally, it reaches $X=0$, where $u\rightarrow\infty$ and $v\rightarrow\infty$, in an infinite affine time according to co-moving observers. This endpoint is also reached in the diagram by the horizon $X=X^{-}$. Similarly, the trajectory of the origin of the radial coordinate ($r=0$) in the FRW interior also terminates there (see Appendix B for the demonstration). Interestingly, the only radial null geodesic that reaches that point, i.e. the final collapsing point, is also a $X=X^{-}$ horizon generator. The rest of radial null geodesics either leave the collapsing dust ball forever at some point or never enter it. We show in Appendix C that the geodesic in question reaches the final point in infinite affine time. The regular quantum gravity core, represented by the timelike line $X=0$, is causally disconnected from the dust content. The diagram may be analytically extended beyond $X=X^{-}$ at the top to reproduce infinitely many copies of the Hayward static geometry there.
IV Discussion
In this work, starting with the Hayward black hole metric and imposing junction conditions on the surface of a collapsing dust ball, we have constructed the complete geometry of a modified OS collapse model, which includes corrections arising from Planck scales. The Hayward exterior is static, whereas the interior is described by a homogeneous cosmology for which we have deduced corrected Friedmann equations. The resulting spacetime is geodesically complete. A power-law contraction of the interior dust matter precedes a smooth transition to a de Sitter phase. Although here the time orientation is chosen so that the model describes gravitational collapse, the cosmological evolution admits also a time-reversed version, in which de Sitter inflation is followed by a a graceful exit to a power-law behavior. Qualitatively, this cosmological evolution resembles the one found in the Starobinsky model [47]. Moreover, cosmic inflation plays in the model an unexpected role, namely, it provides a (quantum) mechanism that decelerates the collapsing matter so that a Schwarzchild-like singularity is never formed.
The conformal diagram of our modified OS collapse model is shown in Fig. 2. This diagram provides the global causal structure of the model. Before computing the diagram, it was not obvious whether the collapsing matter might terminate at the crossing point of the $X^{-}$ horizon and the core, $X=0$. Remarkably, the whole congruence of $r$-constant timelike geodesics actually ends up there forming a caustic. However, the fact that such point is reached after an infinite amount of affine time of the free-falling observers may suggest that the associated divergence of the energy density $\rho\rightarrow\infty$ would be absent in a more realistic collapse scenario accounting for the back-reaction of the Hawking radiation. The diagram representing the evaporation of the Hayward black hole (see Fig. 5 in Ref. [5]) indicates that the black hole lifetime is finite for free-falling observers, so that, in our model, the whole black hole mass could evaporate before $T\rightarrow\infty$. Similarly, the only radial null geodesic reaching the final collapsing point arrives there in infinite affine time. Surprisingly, for an FRW dust cosmology governed by Eqs. (3) and (9), the geodesic completeness could be spoiled. That is, for a cosmological model beyond OS, i.e. $r\in[0,\infty)$, there might be infinitely radial null geodesics reaching $a\rightarrow 0$ in finite affine time. One may speculate that the cosmology studied here is geodesically complete only when surrounded by the static black hole metric. However, the analysis of the corresponding model without the black hole is beyond the scope of this work. Nevertheless, our model teaches us a lesson, namely, that inflationary cosmological spacetimes, where $r\in[0,r_{b}]$, might be past (or future) complete when the exterior vacua are described by smoothly joined black hole metrics. In this way, such models may circumvent the Borde-Guth-Vilenkin theorem [48], which states that spacetimes of this type are incomplete when $r\in[0,\infty)$.
Besides, the diagram indicates that the presence of the inner horizon may probably lead to instabilities, known in the literature as ‘mass inflation’ [49]. In particular, the main reason argued for the appearance of instabilities is that light rays traveling in the vicinity of the inner horizon would experience an arbitrarily high blueshift. However, the overall situation again might be cured by taking into consideration the back-reaction of the Hawking quanta, as advocated by Hayward (since a black hole evaporates, an inner horizon is no more real than the event horizon [5]). Preliminary investigations supporting this possibility were recently conducted in Ref. [17]. Nonetheless, the construction of a conformal diagram that incorporates the consequences of Hawking radiation, at least effectively, is beyond the scope of this work. It would be interesting to see whether Fig. 5 in Ref. [5] is significantly modified when one considers the dust collapse presented here, instead of Vaidya-like ingoing energy flux forming the black hole.
Our model finds motivation in the recently proved fact that the Hayward metric is a vacuum solution of two different effective theories that incorporate quantum gravity corrections, namely, Quasi-Topological gravity [39, 41] and LQG-inspired LTB models [29]. Remarkably, the time dynamics of the scale factor which we have derived (see Fig. 1) are similar to the evolution displayed in Fig. 1 of Ref. [41] and in Fig. 5 of Ref. [29]. This further supports the compatibility of our OS collapse model with those effective approaches.
The considered model presents several clear advantages with respect to other alternatives discussed in the literature. In this sense, it is worth emphasizing the relative simplicity of the cosmological dynamics governed by Eq. (9), the non-singular nature of the resulting spacetime and, most importantly, the inflationary phase that may successfully describe early stages of our universe. In contrast, bouncing OS dust collapse models coming from standard LQG techniques [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] are singular, because they admit Reissner-Nordström-like singularities in vacuo [^5] Notice also that the careful analysis carried out in Refs. [34, 27] shows the singular behavior of the models proposed in Refs. [30, 31, 32, 33] and of their causal structures.. On the other hand, going beyond OS to the case of inhomogeneous dust collapse, the standard LQG techniques lead to shell-crossing singularities [34]. On the contrary, the geometry of the model proposed in this work is smooth everywhere, unlike that in Refs. [43, 38]. Let us also point out that there exist similarities between the dynamics of the model and the quantum-inspired gravitational collapse considered in Ref. [7]. A Taylor expansion of the right-hand side of Eq. (9) around $\rho=0$ yields an infinite tower of (quantum) corrections characterized by consecutive powers of $\rho$ and $l$. These corrections were truncated at $\rho^{3}$ in Ref. [7] (see the discussion around Eq. (23) in that work). However, the cosmological interiors obtained in Ref. [7] may lead to timelike singularities in the exterior, just like in the case of standard LQG models, because the corresponding modified Friedmann equations are similar (see Eq. (24) in Ref. [7] and Eq. (2) in Ref. [25]). The inhomogeneous variant [50] of Ref. [7] is also affected by comparable problems. Thus, we see that a fully detailed study of the causal structure is necessary to decide whether the spacetime is actually non-singular. Indeed, the analysis in this work shows the importance of a rigorous construction of the conformal diagram in collapse models. It would be interesting to see whether some crucial features (for example, singularities) of the models in Refs. [6, 7, 8, 10, 11, 14, 15, 16, 18, 19, 20, 21, 22] remain hidden because their conformal diagrams have not been obtained. Interestingly, our dust collapse model shows also a remarkable resemblance with the model obtained in Ref. [44] using ASG. More concretely, the scale factor and the exterior metric function behave qualitatively in the same way (see Figs. 1 and 2 in that reference), and the dust energy density also diverges. Thus, one can expect that the conformal diagram obtained here (but absent in Ref. [44]) and all the consequences derived from it would apply also to that case. However, our model does not require a running Newton gravitational “constant”, a phenomenon which has not been (at least so far) experimentally supported.
Summarizing, our model represents a clear advance, with important improvements compared to other proposals published in the literature inasmuch as it simultaneously displays the following good physical properties. (i) It does not need to introduce new degrees of freedom, or matter fields additional to those of the case of dust collapse in GR, or a running gravitational constant. (ii) It has a relatively simple cosmological dynamics that may successfully describe the early stages of our universe (with graceful exit from exponential inflation). (iii) The resulting spacetime is nonsingular (this fact was verified by a detailed analysis of the causal structure).
Acknowledgements
Appendix A the Krestchmann scalar for the interior
The Kretschmann scalar for the line element (3), with the scale factor $a$ determined by the inverse of $T(a)$ given in Eq. (7), can be calculated to be
$\begin{split}K_{\mathrm{int}}&=\frac{12\left(a^{2}T^{\prime\prime}(a)^{2}+T^{\prime}(a)^{2}\right)}{a^{4}T^{\prime}(a)^{6}}=\\ &=\frac{12M^{2}\left(5a^{6}r_{b}^{6}+8a^{3}l^{2}Mr_{b}^{3}+32l^{4}M^{2}\right)}{\left(a^{3}r_{b}^{3}+2l^{2}M\right)^{4}}\,.\end{split}$ (13)
Appendix B Construction of the global coordinate system
To cover the exterior region in Fig. 2, we perform the compactification
@@ Appendix B Construction of the global coordinate system
where $c_{v}\in\{0,1\}$, $c_{u}\in\{0,1,2\}$, and $\kappa$ are suitably chosen constants. For the sake of an example, consider the trajectory of the surface given by the pair $\{u(T,R(T)),v(T,R(T)\}$. The $u$ coordinate reaches infinity exactly at $R\rightarrow X^{+}$. So, to obtain a monotonic and compactified coordinate $\tilde{u}$, we have to choose $c_{u}=0$ and the plus sign in front of the $\arctan$ function for $R>X^{+}$, and correspondingly $c_{u}=1$ and the minus sign for $X^{-}<R<X^{+}$. Similarly, the $u$ coordinate reaches minus infinity exactly at $R\rightarrow X^{-}$. Thus, we choose $c_{u}=2$ and the plus sign for $R<X^{-}$. The coordinate $v$ is already monotonic for the trajectory of the surface, so we take $c_{v}=0$ and the plus sign in this case. However, to cover the vacuum part enclosed by $X=X^{-}$ and $X=0$ (see Fig. 2), we have to choose $c_{v}=1$ and the minus sign. Of course, we have to appropriately match signs and constants whenever the surfaces $X^{-}$ or $X^{+}$ are crossed. As for $\kappa$, it is simply a dimensionless, positive constant that allows us to adjust the form of the diagram. The above construction is a simplified (but equally general) version of the procedure presented in Ref. [51].
To extend Eq. (14) to the interior, we follow the method developed in Ref. [46]. We define
@@
$\lim_{T\rightarrow\infty}u_{\mathrm{int}}=\infty,\quad\lim_{T\rightarrow\infty}v_{\mathrm{int}}=\infty.$
@@
We refer to Ref. [46] for the detailed construction and limitations of the presented method.
Appendix C Unique radial null geodesic reaching the final collapsing point
We will now show that the radial null geodesic which is also a $X=X^{-}$ horizon generator (see Fig. 2) takes an infinite affine time to reach the final collapsing point. We rewrite the metric (1) using outgoing Eddington-Finkelstein (EF) coordinates defined by a condition coming from (12), namely $\mathrm{d}u=\mathrm{d}t-\mathrm{d}X/F(X)$. This leads to
@@ Appendix C Unique radial null geodesic reaching the final collapsing point
Let the coordinate $u$ be a parameter of a family of radial null geodesics, so that the tangent vector in EF coordinates is given by $q^{\alpha}=(1,-F(X)/2)$. We have for the four-acceleration $q^{\alpha}\nabla_{\alpha}q^{\beta}=-(\mathrm{d}F/\mathrm{d}X)q^{\beta}/2$. Let $\lambda^{*}$ be the affine parameter. The parameters are related by the differential equation (see Chapter 1.3 of [45] for a general discussion)
@@
where $c_{1}$ and $c_{2}$ are integration constants. For the $X=X^{-}$ horizon generator, we have $\mathrm{d}F/\mathrm{d}X=const.<0$. Thus, when the final collapsing point is reached, with $u\rightarrow\infty$, the affine parameter $\lambda^{*}$ diverges.

### v3

@@ III Conformal Diagram
The integrand in the expression of $v(T,X)$ is regular, whereas the one for $u(T,X)$ is not. However, it can be handled with methods similar to those presented in Ref. Bobula and Pawłowski (2023) (see the discussion around Eqs. 2.29-3.31). These null coordinates still need to be appropriately compactified. To cover the exterior region in the left diagram in Fig. 2, we perform the compactification
@@
where $c_{v}\in\{0,1\}$, $c_{u}\in\{0,1,2\}$, and $\kappa$ are suitably chosen constants. For the sake of an example, consider the trajectory of the surface given by the pair $\{u(T,R(T)),v(T,R(T)\}$. The $u$ coordinate reaches infinity exactly at $R\rightarrow X^{+}$. So, to obtain a monotonic and compactified coordinate $\tilde{u}$, we have to choose $c_{u}=0$ and the plus sign in front of the $\arctan$ function for $R>X^{+}$, and correspondingly $c_{u}=1$ and the minus sign for $X^{-}<R<X^{+}$. Similarly, the $u$ coordinate reaches minus infinity exactly at $R\rightarrow X^{-}$. Thus, we choose $c_{u}=2$ and the plus sign for $R<X^{-}$. The coordinate $v$ is already monotonic for the trajectory of the surface, so we take $c_{v}=0$ and the plus sign in this case. However, to cover the vacuum part enclosed by $X=X^{-}$ and $X=0$ (see Fig. 2), we have to choose $c_{v}=1$ and the minus sign. Of course, we have to appropriately match signs and constants whenever the surfaces $X^{-}$ or $X^{+}$ are crossed. As for $\kappa$, it is simply a dimensionless, positive constant that allows us to adjust the form of the diagram. The above construction is a simplified (but equally general) version of the procedure presented in Ref. Schindler and Aguirre (2018).
To extend Eq. (14) to the interior, we follow the method developed in Ref. Bobula and Pawłowski (2024). We define
@@
$\lim_{T\rightarrow\infty}u_{\mathrm{int}}=\infty,\quad\lim_{T\rightarrow\infty}v_{\mathrm{int}}=\infty\,.$ (17)
@@
We refer to Ref. Bobula and Pawłowski (2024) for the detailed construction and limitations of the presented method for coordinate system extensions.
For the conformal diagram in the right in Fig. 2, being viable for classical OS collapse scenario Oppenheimer and Snyder (1939), we constructed the global coordinate system $(\tilde{u}_{\mathrm{cl}},\tilde{v}_{\mathrm{cl}})$ in an analogous manner as presented in the above paragraphs.
The conformal diagram in the left in Fig. 2 represents the collapse of a dust ball forming a Hayward black hole. The surface of the collapsing ball, at $r=r_{b}$, starts from $X\rightarrow\infty$ and crosses the pair of horizons $X^{+}$ and $X^{-}$, where $u\rightarrow\infty$ and $u\rightarrow-\infty$, respectively. Finally, it reaches $X=0$, where $u\rightarrow\infty$ and $v\rightarrow\infty$, in an infinite affine time according to co-moving observers. This endpoint is also reached in the diagram by the horizon $X=X^{-}$. Similarly, the trajectory of the origin of the radial coordinate ($r=0$) in the FRW interior also terminates there (see the previous paragraphs for the demonstration). Interestingly, the only radial null geodesic that reaches that point, i.e. the final collapsing point, is also a $X=X^{-}$ horizon generator (see Fig. 2). The rest of radial null geodesics either leave the collapsing dust ball forever at some point or never enter it. We will show below that the geodesic in question reaches the final point in infinite affine time.
We rewrite the metric (1) using outgoing Eddington-Finkelstein (EF) coordinates defined by a condition coming from (13), namely $\mathrm{d}u=\mathrm{d}t-\mathrm{d}X/F(X)$. This leads to
@@
Let the coordinate $u$ be a parameter of a family of radial null geodesics, so that the tangent vector in EF coordinates is given by $q^{\alpha}=(1,-F(X)/2)$. We have for the four-acceleration $q^{\alpha}\nabla_{\alpha}q^{\beta}=-(\mathrm{d}F/\mathrm{d}X)q^{\beta}/2$. Let $\lambda^{*}$ be the affine parameter. The parameters are related by the differential equation (see Chapter 1.3 of Poisson (2009) for a general discussion)
@@
where $c_{1}$ and $c_{2}$ are integration constants. For the $X=X^{-}$ horizon generator, we have $\mathrm{d}F/\mathrm{d}X=const.<0$. Thus, when the final collapsing point is reached, with $u\rightarrow\infty$, the affine parameter $\lambda^{*}$ diverges. We have shown that the unique radial null geodesic reaches the final collapsing point in infinite affine time.
We also observe that the regular quantum gravity core, represented by the timelike line $X=0$, is causally disconnected from the dust content. The diagram may be analytically extended beyond $X=X^{-}$ at the top to reproduce infinitely many copies of the Hayward static geometry there.
IV Discussion
In this work, starting with the Hayward black hole metric and imposing junction conditions on the surface of a collapsing dust ball, we have constructed the complete geometry of a modified OS collapse model, which includes corrections arising from Planck scales. The Hayward exterior is static, whereas the interior is described by a homogeneous cosmology for which we have deduced corrected Friedmann equations. The resulting spacetime is geodesically complete. A power-law contraction of the interior dust matter precedes a smooth transition to a de Sitter phase. Although here the time orientation is chosen so that the model describes gravitational collapse, the cosmological evolution admits also a time-reversed version, in which de Sitter inflation is followed by a a graceful exit to a power-law behavior. Qualitatively, this cosmological evolution resembles the one found in the Starobinsky model Starobinsky (1980). Moreover, cosmic inflation plays in the model an unexpected role, namely, it provides a (quantum) mechanism that decelerates the collapsing matter so that a Schwarzchild-like singularity is never formed.
The conformal diagram of our modified OS collapse model is shown in the left in Fig. 2. This diagram provides the global causal structure of the model. Before computing the diagram, it was not obvious whether the collapsing matter might terminate at the crossing point of the $X^{-}$ horizon and the core, $X=0$. Remarkably, the whole congruence of $r$-constant timelike geodesics actually ends up there forming a caustic. However, the fact that such point is reached after an infinite amount of affine time of the free-falling observers may suggest that the associated divergence of the energy density $\rho\rightarrow\infty$ would be absent in a more realistic collapse scenario accounting for the back-reaction of the Hawking radiation. The diagram representing the evaporation of the Hayward black hole (see Fig. 5 in Ref. Hayward (2006)) indicates that the black hole lifetime is finite for free-falling observers, so that, in our model, the whole black hole mass could evaporate before $T\rightarrow\infty$. Similarly, the only radial null geodesic reaching the final collapsing point arrives there in infinite affine time. Surprisingly, an FRW dust cosmology governed by Eqs. (3) and (10) is geodesically incomplete. That is, for a cosmological model beyond OS, i.e. $r\in[0,\infty)$, there will be infinitely many radial null geodesics reaching $a\to 0$ in finite affine time, which implies geodesic incompleteness Borde et al. (2003). In contrast, when the cosmological dust ball is surrounded by a Hayward black hole, geodesic completeness is preserved. The detailed analysis of the corresponding model without the black hole is beyond the scope of this work. Nevertheless, our model teaches us a lesson, namely, that inflationary cosmological spacetimes, where $r\in[0,r_{b}]$, might be past (or future) complete when the exterior vacua are described by smoothly joined black hole metrics. In this way, such models may circumvent the Borde-Guth-Vilenkin theorem Borde et al. (2003), which states that spacetimes of this type are incomplete when $r\in[0,\infty)$.
Besides, the diagram indicates that the presence of the inner horizon may probably lead to instabilities, known in the literature as ‘mass inflation’ Poisson and Israel (1989). In particular, the main reason argued for the appearance of instabilities is that light rays traveling in the vicinity of the inner horizon would experience an arbitrarily high blueshift. However, the overall situation again might be cured by taking into consideration the back-reaction of the Hawking quanta, as advocated by Hayward (since a black hole evaporates, an inner horizon is no more real than the event horizon Hayward (2006)). Preliminary investigations supporting this possibility were recently conducted in Ref. Bonanno et al. (2023). Nonetheless, the construction of a conformal diagram that incorporates the consequences of Hawking radiation, at least effectively, is beyond the scope of this work. It would be interesting to see whether Fig. 5 in Ref. Hayward (2006) is significantly modified when one considers the dust collapse presented here, instead of Vaidya-like ingoing energy flux forming the black hole.
Our model finds motivation in the recently proved fact that the Hayward metric is a vacuum solution of two different effective theories that incorporate quantum gravity corrections, namely, Quasi-Topological gravity Bueno et al. (2025); Arciniega et al. (2020) and LQG-inspired LTB models Giesel et al. (2024c). Remarkably, the time dynamics of the scale factor which we have derived (see Fig. 1) are similar to the evolution displayed in Fig. 1 of Ref. Arciniega et al. (2020) and in Fig. 5 of Ref. Giesel et al. (2024c). This further supports the compatibility of our OS collapse model with those effective approaches. We note that there are prior works on the formation of the Hayward black hole Shojai et al. (2022); Malafarina and Toshmatov (2022). However, in these studies, the Hayward black hole was treated as a non-vacuum solution, leading to critical differences in the physical aspects compared to the model we propose. In Ref. Shojai et al. (2022), a model based on the Einstein Field Equations was derived, wherein the Hayward black hole forms due to the collapse of a polytropic star with non-zero pressure and matter content present both in the interior and the exterior of the star. This model requires exotic matter (negative pressure develops in the star’s interior), and the underlying dynamics differ from those derived in our work for a dust ball interior, as described in Eq. (10). The other study, Ref. Malafarina and Toshmatov (2022), is based on EFEs coupled to nonlinear electrodynamics, where the presence of the electromagnetic field permits the Hayward black hole as a solution. An effective dust collapse model is proposed in this work, where ’semiclassical corrections’ are identified with contributions from nonlinear electrodynamics. The resulting modified Friedmann equation (see Eq. (31) in Ref. Malafarina and Toshmatov (2022)) explicitly depends on the magnetic charge, which is not the case for Eq. (10) in our work. While the trajectories of the collapsing objects discussed in Refs. Shojai et al. (2022); Malafarina and Toshmatov (2022) may be compatible with the geometry we derived, the nonsingular nature of the studied geometry is not evident from these references [^5] A detailed analysis, including the rigorous extraction of the conformal diagram, computation of the Kretschmann scalar, investigation of the final collapsing ’point’, and examination of the only radial null geodesic arriving there, is notably absent in Refs.Shojai et al. (2022); Malafarina and Toshmatov (2022) for the collapse into a Hayward black hole.. Indeed, we believe these studies overlook or underrepresent the most striking feature being present in our model: the role of the cosmic inflation in the collapse setting. The inflation (with the graceful entrance to the de Sitter phase) decelerates the collapsing matter, thereby preventing the formation of a singularity.
The model we have derived presents several clear advantages with respect to other alternatives discussed in the literature. In this sense, it is worth emphasizing the relative simplicity of the cosmological dynamics governed by Eq. (10), the non-singular nature of the resulting spacetime and, most importantly, the inflationary phase that may successfully describe early stages of our universe. In contrast, bouncing OS dust collapse models coming from standard LQG techniques Bobula ; Bobula and Pawłowski (2023); Lewandowski et al. (2023); Giesel et al. (2023, 2024a, 2024b, 2024c); Kelly et al. (2021, 2020); Husain et al. (2022a, b); Fazzini et al. (2024, 2023); Cafaro and Lewandowski (2024); Cipriani et al. (2024); Han et al. (2024) are singular, because they admit Reissner-Nordström-like singularities in vacuo [^6] Notice also that the careful analysis carried out in Refs. Fazzini et al. (2024); Giesel et al. (2024a) shows the singular behavior of the models proposed in Refs. Kelly et al. (2021, 2020); Husain et al. (2022a, b) and of their causal structures.. On the other hand, going beyond OS to the case of inhomogeneous dust collapse, the standard LQG techniques lead to shell-crossing singularities Fazzini et al. (2024). On the contrary, the geometry of the model proposed in this work is smooth everywhere, unlike that in Refs. Han et al. (2023, 2024). Let us also point out that there exist similarities between the dynamics of the model and the quantum-inspired gravitational collapse considered in Ref. Bambi et al. (2013). A Taylor expansion of the right-hand side of Eq. (10) around $\rho=0$ yields an infinite tower of (quantum) corrections characterized by consecutive powers of $\rho$ and $l$. These corrections were truncated at $\rho^{3}$ in Ref. Bambi et al. (2013) (see the discussion around Eq. (23) in that work). However, the cosmological interiors obtained in Ref. Bambi et al. (2013) may lead to timelike singularities in the exterior, just like in the case of standard LQG models, because the corresponding modified Friedmann equations are similar (see Eq. (24) in Ref. Bambi et al. (2013) and Eq. (2) in Ref. Lewandowski et al. (2023)). The inhomogeneous variant Liu et al. (2014) of Ref. Bambi et al. (2013) is also affected by comparable problems. Thus, we see that a fully detailed study of the causal structure is necessary to decide whether the spacetime is actually non-singular. Indeed, the analysis in this work shows the importance of a rigorous construction of the conformal diagram in collapse models. It would be interesting to see whether some crucial features (for example, singularities) of the models in Refs. Barragan et al. (2009); Bambi et al. (2013); Saini and Stojkovic (2014); Frolov (2015); Frolov et al. (2015); Kiefer and Schmitz (2019); Mosani et al. (2020); Góźdź et al. (2023); Shafiee and Bahrampour (2023); Kiefer and Mohaddes (2023); Alonso-Bardaji and Brizuela (2024); Duque (2024); Barca and Montani (2024) remain hidden because their conformal diagrams have not been obtained. Interestingly, our dust collapse model shows also a remarkable resemblance with the model obtained in Ref. Bonanno et al. (2024) using ASG. More concretely, the scale factor and the exterior metric function behave qualitatively in the same way (see Figs. 1 and 2 in that reference), and the dust energy density also diverges. Thus, one can expect that the conformal diagram obtained here (but absent in Ref. Bonanno et al. (2024)) and all the consequences derived from it would apply also to that case. However, our model does not require a running Newton gravitational “constant”, a phenomenon which has not been (at least so far) experimentally supported.
Summarizing, our model represents a clear advance, with important improvements compared to other proposals published in the literature inasmuch as it simultaneously displays the following good physical properties. (i) It does not need to introduce new degrees of freedom, or (exotic) matter fields additional to those of the case of dust collapse in GR, or a running gravitational constant. (ii) It has a relatively simple cosmological dynamics that may successfully describe the early stages of our universe (with graceful exit from exponential inflation). (iii) The resulting spacetime is nonsingular (this fact was verified by a detailed analysis of the causal structure).
Acknowledgements